ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 265, 2025

Argyros, S. A., A. Manoussakis and P. Motakis Variants of the James tree space

1

Arzhakova, E., D. Lind, K. Schmidt and E. Verbitskiy Decimation limits of principal algebraic ℤ^{d}-actions

255

Baake, M., Á. Bustos and A. Nickel Power-free points in quadratic number fields: Stabiliser, dynamics and entropy

79

Bergweiler, W. and A. Tsantaris Hausdorff dimension in quasiregular dynamics

429

Bienvenu, P.-Y. and A. Geroldinger On algebraic properties of power monoids of numerical monoids

867

Bonomo, W., H. Lima and P. Varandas Shape and stability of rotation sets for incompressible flows on the torus

729

Borceux, F., F. Campanini and M. Gran Pretorsion theories in lextensive categories

833

Burniol Clotet, S. Unique ergodicity of horocyclic flows on nonpositively curved surfaces

185

Bustos, Á See M. Baake Campanini, F. See F. Borceux Di Vincenzo, O. M. and V. C. Nardozza Minimal ∗-varieties and

superinvolutions

311

Ding, Y.-H., C.-P. Dong and L. Wei Dirac series of E_{7(7)}  347

Diógenes, R., N. Pinheiro and E. Ribeiro Jr Integral and boundary estimates for critical metrics of the volume functional

901

Dong, C.-P. See Y.-H. Ding

Ercole, G. On a global gradient estimate in p-Laplacian problems  945

Fang, J. See C. Jiang Fehér, L. M. and J. Nagy Additive combinatorics using equivariant

cohomology

533

Galkin, O. E. and I. D. Remizov Upper and lower estimates for rate of convergence in the Chernoff product formula for semigroups of operators

929

Gaster, J. B. and D. T. Wise Bicollapsibility and groups with torsion

501

Geroldinger, A. See P.-Y. Bienvenu Gran, M. See F. Borceux Hillman, J. and D. H. Kochloukova Pro-p completions of P D_{n}-

groups

801

Ioppolo, A. and D. La Mattina Algebras with superautomorphism: simple algebras and codimension growth

479

Ji, K. See C. Jiang Jiang, C., J. Fang and K. Ji Cowen–Douglas operators and the

third of Halmos’ ten problems

379

Kaplan, E. The Magidor iteration and restrictions of ultrapowers to the ground model

605

Karzhemanov, I. and G. Konovalov Rational maps and K3 surfaces

301

Kobayashi, T. and S. Kumashiro The set of trace ideals of curve singularities

153

Kochloukova, D. H. See J. Hillman Konovalov, G. See I. Karzhemanov Kumashiro, S. See T. Kobayashi Li, H. The fundamental groups of presymplectic Hamiltonian G-

manifolds

969

Lima, H. See W. Bonomo Lind, D. See E. Arzhakova Manoussakis, A. See S. A. Argyros La Mattina, D. See A. Ioppolo Mohammadi, A. and N. Tamam Property (τ) in positive character-

istic

569

Monjon, O., J. Scherer and F. Sterck Non-existence of fiberwise localization for crossed modules

217

Motakis, P. See S. A. Argyros Nagy, J. See L. M. Fehér Nardozza, V. C. See O. M. Di Vincenzo Nickel, A. See M. Baake Nyberg-Brodda, C.-F. A word-hyperbolic special monoid with un-

decidable Diophantine problem

723

Peng, Z. and Z. Wang Wavefront sets and descent method for finite symplectic groups

661

Pinheiro, N. See R. Diógenes Polishchuk, A. and C. Wang Linear subspaces in cubic hypersur-

faces

243

Pyörälä, A., P. Shmerkin, V. Suomala and M. Wu Covering the Sierpiński carpet with tubes

769

Remizov, I. D. See O. E. Galkin Ribeiro Jr, E. See R. Diógenes Scherer, J. See O. Monjon Schmidt, K. See E. Arzhakova

Shioya, M. An explicit collapse for a dense filter  467

Shmerkin, P. See A. Pyörälä So, T. and S. Theriault The suspension of a 4-manifold and its

applications

115

Sterck, F. See O. Monjon Suomala, V. See A. Pyörälä Tamam, N. See A. Mohammadi Theriault, S. See T. So Tsantaris, A. See W. Bergweiler Varandas, P. See W. Bonomo Verbitskiy, E. See E. Arzhakova Wang, C. See A. Polishchuk Wang, Z. See Z. Peng Wei, L. See Y.-H. Ding Wise, D. T. See J. B. Gaster Wu, M. See A. Pyörälä